give the number of valance electrons for ICl5
Answer #1

Answer :

In the molecule ICl​​​​5 , I is sp​​​​​3​​​​d hybridised since I make 5 sigma bonds with 5 Cl .

Valance shell electron configuration of I = 5s2 5p5 (7 valance electrons)

Valance shell electron configuration of Cl = 3s2 3p5 (7 valance electrons)

For 5 Cl atom = 5 * 7 = 35 valance electrons

Total valance electrons in ICl​​​​5 = 7 ( from I ) + 35( from Cl) = 42 valance electrons